# matrix,arrows,decorations.pathmorphing

### v2

@@ 2.9. A rigidity criterion
We proceed by induction on $d(S)$. Let us then assume that the statement of the lemma holds for all surfaces of modular dimension $<d(S)$. It is clearly not restrictive to assume that the edge $\{v_{0},v_{1}\}$ in the hypothesis of the theorem belongs to $C_{P}(S)\subset{\check{C}}_{P}(S)$.
By Proposition 2.1, it is enough to prove that the action of the given $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on the set of vertices of the procongruence pants complex ${\check{C}}_{P}(S)$ preserves its set of edges, that is to say, for every edge $\{\alpha_{0},\alpha_{1}\}\in{\check{C}}_{P}(S)_{1}$, there holds $\{f(\alpha_{0}),f(\alpha_{1})\}\in{\check{C}}_{P}(S)_{1}$. By the same argument of the proof of the case $S=S_{0,5}$ of the theorem, it is enough to show that this is the case for a set of representatives of the $\mathrm{P}{\check{\Gamma}}(S)$-orbits in ${\check{C}}_{P}(S)_{1}$. In particular, we can assume that $\{\alpha_{0},\alpha_{1}\}\in C_{P}(S)_{1}\subset{\check{C}}_{P}(S)_{1}$ as well.
For $d(S)>2$, the complexes $C_{b}(S)$ (for $g(S)=0$), $C_{0b}(S)$ (for $g(S)=1$) and $C_{0}(S)$ (for $g(S)>1$) are connected. This implies that there is a set ${\gamma}_{1},\dots,{\gamma}_{k}$ of simple closed curves on $S$, where ${\gamma}_{i}$, for $i=1,\dots,k$, is either nonseparating or bounds a $2$-punctured disc, such that the edge $\{v_{0},v_{1}\}$ is contained in $L_{{\gamma}_{1}}$, the edge $\{\alpha_{0},\alpha_{1}\}$ is contained in $L_{{\gamma}_{k}}$ and the intersection $L_{{\gamma}_{i}}\cap L_{{\gamma}_{i+1}}$, for $11\leq i\leq k-$ 1, contains at least an edge of ${\check{C}}_{P}(S)$.
@@
After composing $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ with an element in the image of $\operatorname{Inn}({\check{\Gamma}}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$, we can assume that $f$ preserves the procyclic subgroup ${\hat{\mathrm{I}}}_{{\gamma}_{i}}$ and then acts on the vertex set of the subgraph $L_{{\gamma}_{i}}$, which identifies with the set of $(d(S)-1)$-simplices of the star of ${\gamma}_{i}$ in ${\check{C}}(S)$.
@@ 3.1. Centralizers of antiholomorphic involutions
The centralizer of $\iota$ in ${\Gamma}^{\pm}(S)$ has a simple description. Let $S_{\iota}:=S/\langle\iota\rangle$ be the quotient surface. Let $\operatorname{Fix}(\iota)$ be the fixed point set of $\iota$. Then, $\operatorname{Fix}(\iota)$ is the union of a (possibly empty) set of disjoint simple closed curves on $S$ and the quotient surface $S_{\iota}$ is orientable if and only if $S\smallsetminus\operatorname{Fix}(\iota)$ is not connected. Moreover, if $\operatorname{Fix}(\iota)\neq\emptyset$, then $S_{\iota}$ is a surface with boundary $\partial S_{\iota}$, which coincides with the image of $\operatorname{Fix}(\iota)$ in $S_{\iota}$ (cf. Proposition 1.2 in [24]). Let us denote by $\operatorname{Map}(S_{\iota})$ the group of isotopy classes of self-diffeomorphisms of the (possibly non-orientable) surface $S_{\iota}$. We then have:
@@
$1\to\langle\iota\rangle\to Z_{{\Gamma}^{\pm}(S)}(\iota)\to\operatorname{Map}(S_{\iota})\to 1$
@@
Let us then assume that $\operatorname{Fix}(\iota)\neq\emptyset$ and $S\smallsetminus\operatorname{Fix}(\iota)$ is connected. The surface $S\smallsetminus\operatorname{Fix}(\iota)$ identifies with the orientation cover of $S_{\iota}\smallsetminus\partial S_{\iota}$, so that every self-homeomorphism of $S_{\iota}\smallsetminus\partial S_{\iota}$ lifts to $S\smallsetminus\operatorname{Fix}(\iota)$. Since every self-homeomorphism of $S$ which commutes with $\iota$ preserves the fixed point set $\operatorname{Fix}(\iota)$, the conclusion follows.
@@
3.2. The fixed point set of an antiholomorphic involution in the augmented Teichmüller space
Let ${\mathcal{T}}(S)$ be the Teichmüller space associated to the surface $S$ endowed with the Weil-Petersson metric. From Lemma 3.5 in [22], it follows that the fixed point set ${\mathcal{T}}(S)^{\iota}$ of an antiholomorphic involution $\iota\in{\Gamma}^{\pm}(S)$ is a nonempty and connected real submanifold of ${\mathcal{T}}(S)$ of (real) dimension $d(S)$ (cf. Corollary 3.8 in [22]).
Since the augmented Teichmüller space $\overline{{\mathcal{T}}}(S)$ is the completion of the Teichmüller space ${\mathcal{T}}(S)$ with respect to the Weil-Petersson metric (cf. Section 2.1), from Theorem 5 in [26], it follows that the fixed point set $\overline{{\mathcal{T}}}(S)^{\iota}$ coincides with the closure of the fixed point set ${\mathcal{T}}(S)^{\iota}$ in $\overline{{\mathcal{T}}}(S)$.
Theorem 5 in [26] and Lemma 3.5 in [22] then imply that, for ${\sigma}\in C(S)_{k}$, the fixed point set $\partial\overline{{\mathcal{T}}}(S)_{\sigma}^{\iota}$ of the corresponding closed stratum of $\partial\overline{{\mathcal{T}}}(S)$ is nonempty if and only if ${\sigma}\in C(S)_{k}^{\iota}$ and that, for all ${\sigma}\in C(S)_{k}^{\iota}$, there holds $\partial\overline{{\mathcal{T}}}(S)_{\sigma}^{\iota}\cong\overline{{\mathcal{T}}}(S\smallsetminus{\sigma})^{\iota}$. We sum up the above discussion in the following proposition:
@@ 3.2. The fixed point set of an antiholomorphic involution in the augmented Teichmüller space
The closed strata of codimension $k+1$ in the boundary of the fixed point locus $\overline{{\mathcal{T}}}(S)^{\iota}$ are parameterized by the fixed point set $C(S)_{k}^{\iota}$, for $k\geq 0$.
@@
Note that, for all $k\geq 0$, the action of the centralizer $Z_{{\Gamma}^{\pm}(S)}(\iota)$ on $C(S)_{k}$ preserves the fixed point set $C(S)_{k}^{\iota}$ and acts on the latter with a finite number of orbits.
@@ 3.3. Antiholomorphic involutions of the profinite mapping class group
Let us now assume that $g(S)\leq 2$. By the congruence subgroup property in genus $\leq 2$, we then have ${\check{\Gamma}}(S)\cong{\widehat{\Gamma}}(S)$ and so ${\check{\Gamma}}^{\pm}(S)\cong{\widehat{\Gamma}}^{\pm}(S)$. The augmentation map ${\Gamma}^{\pm}(S)\to\mathbb{Z}/2$ induces an augmentation map ${\widehat{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2$ and we define an antiholomorphic involution of ${\widehat{\Gamma}}^{\pm}(S)$ to be an element of order $2$ whose image by the augmentation map is nontrivial.
It is well known that ${\Gamma}(S)$ is a good group for $g(S)\leq 2$. The same then holds for ${\Gamma}^{\pm}(S)$ and, from Corollary 3.5 in [21], it follows that the natural inclusion ${\Gamma}^{\pm}(S)<{\widehat{\Gamma}}^{\pm}(S)$ induces a bijection between the conjugacy classes of elements of prime order in ${\Gamma}^{\pm}(S)$ and in ${\widehat{\Gamma}}^{\pm}(S)$. In particular, we have:
@@
For $g(S)\leq 2$, every antiholomorphic involution of ${\widehat{\Gamma}}^{\pm}(S)$ is conjugated to an antiholomorphic involution inside ${\Gamma}^{\pm}(S)$.
3.4. The fixed point set of an antiholomorphic involution in the procongruence moduli stack
A description of the fixed point locus $\overline{{\mathbb{M}}}(S)^{\iota}$ for the action of $\iota$ on $\overline{{\mathbb{M}}}(S)$ for $g(S)\leq 2$ is implicit in the proof of Corollary 3.15 in [22] where it is also proved that the centralizer $Z_{{\widehat{\Gamma}}^{\pm}(S)}(\iota)$ coincides with the closure of $Z_{{\Gamma}^{\pm}(S)}(\iota)$ in ${\widehat{\Gamma}}^{\pm}(S)$. Here, we prefer to give an alternative proof of these statements since the argument can be substantially simplified using Corollary 3.5 in [21]:
@@ 3.4. The fixed point set of an antiholomorphic involution in the procongruence moduli stack
For $g(S)\leq 2$, the fixed point set locus $\overline{{\mathbb{M}}}(S)^{\iota}$ of an antiholomorphic involution $\iota\in{\Gamma}^{\pm}(S)<{\widehat{\Gamma}}^{\pm}(S)$ is smooth, irreducible and contains ${{\mathbb{M}}}(S)^{\iota}$ as an open dense subspace. The centralizer $Z_{{\widehat{\Gamma}}^{\pm}(S)}(\iota)$ coincides with the closure of $Z_{{\Gamma}^{\pm}(S)}(\iota)$ in ${\widehat{\Gamma}}^{\pm}(S)$.
@@
For every torsion free characteristic level ${\Gamma}^{\lambda}$ of ${\Gamma}(S)$, let $({\mathcal{M}}(S)^{\lambda},\iota)$ be the real complex manifold with equivariant fundamental group (cf. Section 3 in [12]) isomorphic to ${\Gamma}^{\lambda}\cdot\langle\iota\rangle$. By Theorem 3.6 in [22], there is a natural bijection between the set of connected components of the real locus of $({\mathcal{M}}(S)^{\lambda},\iota)$ and conjugacy classes of involutions in ${\Gamma}^{\lambda}\cdot\langle\iota\rangle$. For $g(S)\leq 2$, the latter is a good group and then, by Corollary 3.5 in [21], we conclude that there is a bijective correspondence between the set of connected components of the real locus of $({\mathcal{M}}(S)^{\lambda},\iota)$ and conjugacy classes of involutions in ${\widehat{\Gamma}}^{\lambda}\cdot\langle\iota\rangle$. Passing to the inverse limit over all such level structures, since $\cap_{{\lambda}\in\Lambda}{\widehat{\Gamma}}^{\lambda}\cdot\langle\iota\rangle=\langle\iota\rangle$, it follows that the fixed point set locus ${\mathbb{M}}(S)^{\iota}$ is connected, smooth (and so irreducible) and that the centralizer $Z_{{\widehat{\Gamma}}^{\pm}(S)}(\iota)$ identifies with the covering transformation group of the cover from ${\mathbb{M}}(S)^{\iota}$ to the connected component of ${\mathcal{M}}(S)_{\mathbb{R}}$ associated to the conjugacy class of $\iota$ in ${\widehat{\Gamma}}^{\pm}(S)$.
The statements in the lemma about the fixed point set locus $\overline{{\mathbb{M}}}(S)^{\iota}$ then follow from the fact that this locus is the closure of ${\mathbb{M}}(S)^{\iota}$ in $\overline{{\mathbb{M}}}(S)$. ∎
The closed boundary strata of $\overline{{\mathbb{M}}}(S)$ are parameterized by the simplices of the procongruence curve complex ${\check{C}}(S)$ and, for ${\sigma}\in C(S)\subset{\check{C}}(S)$, there holds $\partial\overline{{\mathbb{M}}}(S)_{\sigma}\cong\overline{{\mathbb{M}}}(S\smallsetminus{\sigma})$, where we denote by $\partial\overline{{\mathbb{M}}}(S)_{\sigma}$ the closed stratum of $\partial\overline{{\mathbb{M}}}(S)=\overline{{\mathbb{M}}}(S)\smallsetminus{\mathbb{M}}(S)$ parameterized by a simplex ${\sigma}\in{\check{C}}(S)$. From Proposition 3.2 and Proposition 3.5, it then follows:

### v3

@@ 2.9. A rigidity criterion
We proceed by induction on $d(S)$. Let us then assume that the statement of the lemma holds for all surfaces of modular dimension <d(S)<d(S). By the isomorphism (8), it is enough to prove that the action of the given automorphism $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on the set of vertices of the procongruence pants complex ${\check{C}}_{P}(S)$ preserves its set of edges, that is to say, for every edge $\{\alpha_{0},\alpha_{1}\}\in{\check{C}}_{P}(S)_{1}$, there holds $\{f(\alpha_{0}),f(\alpha_{1})\}\in{\check{C}}_{P}(S)_{1}$.
Let us show that, by an argument similar to the proof of the case $S=S_{0,5}$ of the theorem, it is actually enough to show that this is the case for a set of representatives of the $\mathrm{P}{\check{\Gamma}}(S)$-orbits in ${\check{C}}_{P}(S)_{1}$. Thus, assuming that $e=\{w_{0},w_{1}\}\in{\check{C}}_{P}(S)_{1}$ is such that $f(e)=\{f(w_{0}),f(w_{1})\}\in{\check{C}}_{P}(S)_{1}$, let us show that, if $e^{\prime}=\{\alpha_{0},\alpha_{1}\}\in{\check{C}}_{P}(S)_{1}$ has the property that there is an element $x\in{\check{\Gamma}}(S)$ such that $\alpha_{i}=x\cdot w_{i}\cdot x^{-1}$, for $i=0,1$, then we also have $f(e^{\prime})=\{f(\alpha_{0}),f(\alpha_{1})\}\in{\check{C}}_{P}(S)_{1}$. Now, for $i=0,1$, we have the identities:
$f(\alpha_{i})=f(x\cdot w_{i}\cdot x^{-1})=f(x)\cdot f(w_{i})\cdot f(x)^{-1},$
so that $\{f(\alpha_{0}),f(\alpha_{1})\}=\operatorname{inn}f(x)(\{f(w_{0}),f(w_{1})\})\in{\check{C}}_{P}(S)_{1}$.
By the remark above, in particular, we can assume that $\{\alpha_{0},\alpha_{1}\}\in C_{P}(S)_{1}\subset{\check{C}}_{P}(S)_{1}$.
For $d(S)>2$, the complexes $C_{b}(S)$ (for $g(S)=0$), $C_{0b}(S)$ (for $g(S)=1$) and $C_{0}(S)$ (for $g(S)>1$) are connected. This implies that there is a set ${\gamma}_{1},\dots,{\gamma}_{k}$ of simple closed curves on $S$, where ${\gamma}_{i}$, for $i=1,\dots,k$, is either nonseparating or bounds a $2$-punctured disc, such that the edge $\{v_{0},v_{1}\}$ is contained in $L_{{\gamma}_{1}}$, the edge $\{\alpha_{0},\alpha_{1}\}$ is contained in $L_{{\gamma}_{k}}$ and the intersection $L_{{\gamma}_{i}}\cap L_{{\gamma}_{i+1}}$, for $1\leq i\leq k-1$, contains at least an edge of ${\check{C}}_{P}(S)$.
@@
After composing $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ with an element in the image of $\operatorname{Inn}({\check{\Gamma}}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$, we can assume that $f$ preserves the procyclic subgroup ${\hat{\mathrm{I}}}_{{\gamma}_{i}}$ and then acts on the vertex set of the subgraph $L_{{\gamma}_{i}}$, which identifies with the set of (d(S)1)(d(S)-1)-simplices of the star of ${\gamma}_{i}$ in ${\check{C}}(S)$.
@@ 3.1. Centralizers of antiholomorphic involutions
The centralizer of $\iota$ in ${\Gamma}^{\pm}(S)$ has a simple description. Let $S_{\iota}:=S/\langle\iota\rangle$ be the quotient surface. Let $\operatorname{Fix}(\iota)$ be the fixed-point set of $\iota$. Then, $\operatorname{Fix}(\iota)$ is the union of a (possibly empty) set of disjoint simple closed curves on $S$ and the quotient surface $S_{\iota}$ is orientable if and only if $S\smallsetminus\operatorname{Fix}(\iota)$ is not connected. Moreover, if $\operatorname{Fix}(\iota)\neq\emptyset$, then $S_{\iota}$ is a surface with boundary $\partial S_{\iota}$, which coincides with the image of $\operatorname{Fix}(\iota)$ in $S_{\iota}$ (cf. [39, Proposition 1.2]). Let us denote by $\operatorname{Map}(S_{\iota})$ the group of isotopy classes of self-diffeomorphisms of the (possibly non-orientable) surface $S_{\iota}$. We then have:
@@
$1\to\langle\iota\rangle\to Z_{{\Gamma}^{\pm}(S)}(\iota)\to\operatorname{Map}(S_{\iota})\to 1.$
@@
Let us then assume that $\operatorname{Fix}(\iota)\neq\emptyset$ and $S\smallsetminus\operatorname{Fix}(\iota)$ is connected. The surface $S\smallsetminus\operatorname{Fix}(\iota)$ identifies with the orientation cover of $S_{\iota}\smallsetminus\partial S_{\iota}$, so that every self-homeomorphism of $S_{\iota}\smallsetminus\partial S_{\iota}$ lifts to $S\smallsetminus\operatorname{Fix}(\iota)$. Since every self-homeomorphism of $S$ which commutes with $\iota$ preserves the fixed-point set $\operatorname{Fix}(\iota)$, the conclusion follows.
@@
3.2. The fixed-point set of an antiholomorphic involution in the augmented Teichmüller space
As it is customary, for a space $X$ endowed with an involution $\iota$, we denote by $X^{\iota}$ its fixed-point set. Let then ${\mathcal{T}}(S)$ be the Teichmüller space associated to the surface $S$ endowed with the Weil–Petersson metric. From [33, Lemma 3.5], it follows that the fixed-point set ${\mathcal{T}}(S)^{\iota}$ of an antiholomorphic involution $\iota\in{\Gamma}^{\pm}(S)$ is a nonempty and connected real analytic submanifold of ${\mathcal{T}}(S)$ of (real) dimension $d(S)$. We have:
Lemma 3.2.
The fixed-point set $\overline{{\mathcal{T}}}(S)^{\iota}$ for the action of $\iota$ on the augmented Teichmüller space $\overline{{\mathcal{T}}}(S)$ is the closure of the fixed-point set ${\mathcal{T}}(S)^{\iota}$ in $\overline{{\mathcal{T}}}(S)$.
Proof.
The augmented Teichmüller space $\overline{{\mathcal{T}}}(S)$ is the completion of the Teichmüller space ${\mathcal{T}}(S)$ with respect to the Weil–Petersson metric (cf. Section 2.1). The conclusion then follows from the fact that, for $x\in{\mathcal{T}}(S)^{\iota}$ and $y\in\partial\overline{{\mathcal{T}}}(S)^{\iota}:=\overline{{\mathcal{T}}}(S)^{\iota}\smallsetminus{\mathcal{T}}(S)^{\iota}$, the unique geodesic connecting $x$ and $y$ (cf. the discussion preceding [42, Theorem 5]) is contained in $\overline{{\mathcal{T}}}(S)^{\iota}$ and intersects ${\mathcal{T}}(S)^{\iota}$ in an open dense subset (cf. [42, Theorem 5]). ∎
For an antiholomorphic involution $\iota$, let us denote by $C(S)^{\iota}_{k}$ the set of $k$-simplices of the abstract simplicial complex $C(S)$ fixed by $\iota$. By [42, Theorem 5] and [33, Lemma 3.5], we then have that, for ${\sigma}\in C(S)_{k}$, the fixed-point set $\partial\overline{{\mathcal{T}}}(S)_{\sigma}^{\iota}$ of the corresponding closed stratum of $\partial\overline{{\mathcal{T}}}(S)$ is nonempty if and only if ${\sigma}\in C(S)_{k}^{\iota}$. Moreover, by Lemma 3.2, for all ${\sigma}\in C(S)_{k}^{\iota}$, there holds $\partial\overline{{\mathcal{T}}}(S)_{\sigma}^{\iota}\cong\overline{{\mathcal{T}}}(S\smallsetminus{\sigma})^{\iota}=\overline{{\mathcal{T}}(S\smallsetminus{\sigma})^{\iota}}$. Let us sum up the above discussion in the following proposition:
@@ 3.2. The fixed-point set of an antiholomorphic involution in the augmented Teichmüller space
The closed irreducible strata of codimension $k+1$ in the boundary of the fixed-point locus $\overline{{\mathcal{T}}}(S)^{\iota}$ are parameterized by the fixed-point set $C(S)_{k}^{\iota}$, for $k\geq 0$.
@@
Note that, for all $k\geq 0$, the action of the centralizer $Z_{{\Gamma}^{\pm}(S)}(\iota)$ on $C(S)_{k}$ preserves the fixed-point set $C(S)_{k}^{\iota}$ and acts on the latter with a finite number of orbits.
@@ 3.3. Antiholomorphic involutions of the profinite mapping class group
Let us assume that $g(S)\leq 2$. By the congruence subgroup property in genus $\leq 2$, we then have that $\mathrm{P}{\check{\Gamma}}(S)\cong\mathrm{P}{\widehat{\Gamma}}(S)$ and so $\mathrm{P}{\check{\Gamma}}(S)^{\pm}\cong\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ and ${\check{\Gamma}}^{\pm}(S)\cong{\widehat{\Gamma}}^{\pm}(S)$. The augmentation map ${\Gamma}^{\pm}(S)\to\mathbb{Z}/2$ induces an augmentation map ${\widehat{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2$ and we define an antiholomorphic involution of ${\widehat{\Gamma}}^{\pm}(S)$ to be an element of order $2$ whose image by the augmentation map is nontrivial.
@@
For $g(S)\leq 2$, we have:
  1. (i)
    ${\Gamma}^{\pm}(S)$ is a good group, that is to say the natural homomorphism ${\Gamma}^{\pm}(S)\hookrightarrow{\widehat{\Gamma}}^{\pm}(S)$ induces an isomorphism on (continuous) cohomology with finite coefficients.
  2. (ii)
    The natural homomorphism ${\Gamma}^{\pm}(S)\hookrightarrow{\widehat{\Gamma}}^{\pm}(S)$ induces a bijection between the sets of conjugacy classes of involutions in the two groups.
  3. (iii)
    For every involution $\iota\in{\Gamma}^{\pm}(S)\subset{\widehat{\Gamma}}^{\pm}(S)$, the centralizer $Z_{{\widehat{\Gamma}}^{\pm}(S)}(\iota)$ coincides with the closure of $Z_{{\Gamma}^{\pm}(S)}(\iota)$ in ${\widehat{\Gamma}}^{\pm}(S)$.
The same statement holds if we replace ${\Gamma}^{\pm}(S)$ by $\mathrm{P}\Gamma^{\pm}(S)$.
Proof.
(i): It is well known and easy to prove that $\mathrm{P}\Gamma(S)$ is a good group for $g(S)\leq 2$ (cf. for instance, the closing remarks in [37] or [4, Proposition 8.5]). The conclusion then follows from the Hochschild–Lyndon–Serre spectral sequence of a group extension.
(ii) and (iii): These follow from the first item and [10, Corollary B]. ∎
3.4. The center of the extended procongruence mapping class group
The following result is an almost immediate consequence of [9, Theorem 4.14]:
@@ 3.4. The center of the extended procongruence mapping class group
Let $U$ be an open subgroup of ${\check{\Gamma}}^{\pm}(S)$.
  1. (i)
    For $S\neq S_{0,4},S_{1,1},S_{1,2}$ and $S_{2}$, the centralizer of $U$ in ${\check{\Gamma}}^{\pm}(S)$ is trivial.
  2. (ii)
    For $S=S_{1,1},S_{1,2}$ and $S_{2}$, the centralizer of $U$ in ${\check{\Gamma}}^{\pm}(S)$ is generated by the hyperelliptic involution $\upsilon$}.
  3. (iii)
    For $S=S_{0,4}$, we have $Z({\check{\Gamma}}^{\pm}(S))=\{1\}$ and if $U\subseteq\mathrm{P}{\check{\Gamma}}^{\pm}(S)$, then $Z_{{\check{\Gamma}}^{\pm}(S)}(U)=K_{4}$, where $K_{4}\cong\{\pm 1\}\times\{\pm 1\}$ is the Klein subgroup of ${\Gamma}_{0,[4]}$.
@@
An antiholomorphic involution $\iota\in{\Gamma}^{\pm}(S)$ acts by conjugation on Dehn twist powers $\tau_{\gamma}^{k}\in\mathrm{P}\Gamma(S)$, for $k\in\mathbb{N}^{+}$, by the formula $\iota\tau_{\gamma}^{k}\iota=\tau_{\iota({\gamma})}^{-k}$. This implies that the outer representation associated to the short exact sequence $1\to{\check{\Gamma}}(S)\to{\check{\Gamma}}(S)^{\pm}\to\mathbb{Z}/2\to 1$ is nontrivial when restricted to the subgroup $U$. Together with [9, Theorem 4.14], this implies all claims of the proposition. ∎
3.5. The fixed-point set of an antiholomorphic involution in the procongruence moduli stack
The main result of this section is Proposition 3.9, which is a key ingredient for the proof of the genus zero case of Theorem 1.4 (cf. Section 4.3).
For every torsion-free characteristic level ${\Gamma}^{\lambda}$ of ${\Gamma}(S)$ and an antiholomorphic involution $\iota\in{\Gamma}^{\pm}(S)$, let $({\mathcal{M}}(S)^{\lambda},\iota)$ be the real complex manifold with equivariant fundamental group isomorphic to ${\Gamma}^{\lambda}\cdot\langle\iota\rangle$ (cf. [19, Section 3]). Let ${\overline{\mathcal{M}}}(S)^{\lambda}$ be the DM compactification of ${\mathcal{M}}(S)^{\lambda}$. The action of $\iota$ on ${\mathcal{M}}(S)^{\lambda}$ extends to ${\overline{\mathcal{M}}}(S)^{\lambda}$, which then acquires a structure of real complex analytic variety.
Let us recall that we defined $\overline{{\mathbb{M}}}(S)$ to be the inverse limit of all compactified geometric level structure over ${\overline{\mathcal{M}}}(S)$ and that there is a natural embedding $\overline{{\mathcal{T}}}(S)\hookrightarrow\overline{{\mathbb{M}}}(S)$ (cf. Section 2). Therefore, the action of $\iota$ on the augmented Teichmüller space $\overline{{\mathcal{T}}}(S)$ extends to $\overline{{\mathbb{M}}}(S)$ (and, in particular, to ${\mathbb{M}}(S)$). We have:
Lemma 3.7.
The fixed-point set $\overline{{\mathbb{M}}}(S)^{\iota}$ is the inverse limit of an inverse system of real analytic manifolds of dimension $d(S)$ and contains ${{\mathbb{M}}}(S)^{\iota}$ as an open dense subspace.
Proof.
For the first claim, by the definitions involved, it is enough to prove that, for a cofinal system of congruence levels $\{{\Gamma}^{\lambda}\}_{{\lambda}\in{\Lambda}^{\prime}}$ of ${\Gamma}(S)$, the fixed-point set $({\overline{\mathcal{M}}}(S)^{\lambda})^{\iota}$ is smooth of real dimension $d(S)$. From [7, Theorem 3.11], it follows that, at least, there is such a cofinal system with the property that the associated compactified level structures ${\overline{\mathcal{M}}}(S)^{\lambda}$ are smooth complex manifolds for all ${\lambda}\in{\Lambda}^{\prime}$. Since, as we observed above, the fixed-point set ${\mathcal{T}}(S)^{\iota}$ is nonempty, we also know that the fixed-point set $({\overline{\mathcal{M}}}(S)^{\lambda})^{\iota}$ is nonempty. This implies (cf. [33, Section 3.1]) that the fixed-point set $({\overline{\mathcal{M}}}(S)^{\lambda})^{\iota}$ is a smooth real analytic subspace of ${\overline{\mathcal{M}}}(S)^{\lambda}$ of (real) dimension $d(S)$, for all ${\lambda}\in{\Lambda}^{\prime}$. The claim then follows.
For the second claim, it is enough to prove that, for every characteristic level ${\Gamma}^{\lambda}$ of ${\Gamma}(S)$, the fixed-point set $({\mathcal{M}}(S)^{\lambda})^{\iota}$ is dense in the fixed-point set $({\overline{\mathcal{M}}}(S)^{\lambda})^{\iota}$. Every irreducible component (as a real analytic variety) of the fixed-point set $({\overline{\mathcal{M}}}(S)^{\lambda})^{\iota}$ is the image of the fixed-point set $\overline{{\mathcal{T}}}(S)^{\iota^{\prime}}$ in the quotient ${\overline{\mathcal{M}}}(S)^{\lambda}=\overline{{\mathcal{T}}}(S)/{\Gamma}^{\lambda}$, for some antiholomorphic involution $\iota^{\prime}$ such that $\iota^{\prime}\equiv\iota\mod{\Gamma}^{\lambda}$. Hence, the conclusion follows from Lemma 3.2. ∎
We then have: